Bylarus:

OpenStudy (anonymous): Lines CD and DE are tangent to circle A as shown below: Lines CD and DE are tangent to circle A and intersect at point D. Arc CE measures 130 degrees. Point B lies on circle A. If arc CE is 130°, what is the measure of ∠CDE? 20° 40° 42.5° 50°

jhonyy9:

do you can post an image about this ?

Bylarus:

yea

jhonyy9:

do you know how many degrees has a circle ?

Bylarus:

180 I think or 360

Bylarus:

or it has none?

Bylarus:

the degree in thepicture i sent says 130

jhonyy9:

a circle has 360 degrees

Bylarus:

yea

jhonyy9:

you know arc CE equal 130 degrees can you calcule arc CBE howmany degrees has ?

Bylarus:

i don't know how unfortunately

Bylarus:

I think they are all the same because of the distance between each other

Bylarus:

so 130 degrees

Bylarus:

?

jhonyy9:

remember a circle has 360 degrees arc CE has 130 degrees how you get the measure of arc CBE ?

Bylarus:

divide 360 by 130

jhonyy9:

not really

jhonyy9:

use a+b = 360 a=130 how you calcule the value of b ?

Bylarus:

i got 230

Bylarus:

thats b

Bylarus:

230 right?

jhonyy9:

yes ofc so you get arc CBE has 230 degrees

Bylarus:

yea

Bylarus:

now what next

jhonyy9:

do you know this formula ? exterior angle measure = (big arc measure - small arc measure)/2

jhonyy9:

or look please on this website https://www.mathwarehouse.com/geometry/circle/tangents-secants-arcs-angles.php

mhanifa:

You can use a simple rule: The angle between two tangents is supplementary with minor arc.

mhanifa:

It means mCE + mCDE = 180

jhonyy9:

@extrinix

Extrinix:

@jhonyy9 wrote:
do you know this formula ? exterior angle measure = (big arc measure - small arc measure)/2
@mhanifa wrote:
It means mCE + mCDE = 180
You guys would both be correct, jhonyy9: \(m\angle D = \dfrac{230^\circ -130^\circ}{2}\) simplify the top \(m\angle D = \dfrac{100}{2}^\circ \rightarrow 50^\circ\) convert, \(m\angle D\) is \(50^\circ\) mhanifa: \(180^\circ = m\angle CDE + 130^\circ\) subtract 130 from both sides \(50^\circ = m\angle CDE\) \(m\angle D\) is \(50^\circ\)